OpenStudy (star):

(question will be rewritten with proper equations) determine the convergence/divergence of the series sum from 2 to infinity of An whose n-th term is given by n^2/n!

OpenStudy (star):

determine the convergence/divergence of \[\sum_{n=2}^{\infty} a _{n}\] whose n-th term is given by \[a _{n} = n^2/n!\]